OpenStudy (anonymous):

Which of the following is the rational exponent expression of sixth root of f?

OpenStudy (jdoe0001):

easy, is "blank", since the choices are "blank"

OpenStudy (jdoe0001):

keep in mind that \(\Large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\)